OpenStudy (anonymous):

Limits The value of f(0) so that the f(x) is continuous at x=0.

OpenStudy (dls):

and f(x)=?

OpenStudy (anonymous):

\[\Large f(x)=\frac{(27-2x)^{1/3}-3}{9-3(243+5x)^{1/5}}\]

OpenStudy (dls):

L hospital B)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

trying

OpenStudy (dls):

\[\Large f'(x)=\frac{\frac{1}{3}(27-2x)^\frac{-2}{3}\times -2}{-3(\frac{1}{5}(243+5x)^\frac{-4}{5}\times 5} \]

OpenStudy (dls):

try subbing x=0 now

OpenStudy (anonymous):

i am getting 37.44

OpenStudy (anonymous):

and ans. is 2

OpenStudy (dls):

im getting 2 only,recheck ur calculation

OpenStudy (anonymous):

ok

OpenStudy (dls):

\[\Huge \frac{\frac{-27}{2}}{\frac{1}{-27 \times 5} \times 5}\]

OpenStudy (dls):

sorry its 2/27 in numerator,typo*